# Equivalence of scalar-tensor theories and scale-dependent gravity

### v1

@@ I Introduction
Despite the overwhelming success of general relativity (GR), evidence has accumulated over the past century that motivates searches for alternative or modified theories of gravity. In particular, questions about the nature of dark energy and dark matter remain among the most significant problems in gravitational physics and cosmology. While it is possible to account for dark energy within GR in the form of a cosmological constant, this leads to severe fine-tuning problems Weinberg (1989). In addition, despite substantial theoretical and experimental efforts, the search for direct evidence of particle dark matter has so far been without success. Consequently, modified theories of gravity are extensively discussed in the literature and many different models have emerged Clifton et al. (2012). Studying the similarities, differences and interrelations of the various approaches is essential for our understanding of gravity as a whole.
@@
A simple way of modifying GR is by coupling an additional scalar field to the metric tensor, resulting in a so-called scalar-tensor theory (STT) of gravity Fujii and Maeda (2007). This can be motivated, for example, by higher-dimensional theories whose compactified extra dimensions give rise to additional scalar degrees of freedom Wehus and Ravndal (2004). Due to the coupling of a scalar field to the metric tensor, we would expect to see an additional fifth force of Nature. However, fifth forces are already heavily constrained by Solar System-based tests Dickey et al. (1994); Adelberger et al. (2003); Kapner et al. (2007). In order to circumvent such experimental bounds, some STTs are equipped with a screening mechanism that renders the fifth force feeble in environments of sufficiently large mass densities. Amongst the most prominent screened scalar field models are the chameleon Khoury and Weltman (2004a, b), the symmetron Dehnen et al. (1992); Gessner (1992); Damour and Polyakov (1994); Pietroni (2005); Olive and Pospelov (2008); Brax et al. (2010); Hinterbichler and Khoury (2010); Hinterbichler et al. (2011), and the environment-dependent dilaton Damour and Polyakov (1994); Gasperini et al. (2002); Damour et al. (2002a, b); Brax et al. (2010, 2011, 2022). The first two have recently been proposed as alternatives to particle dark matter Burrage et al. (2017); O’Hare and Burrage (2018); Burrage et al. (2019a); Käding (2023); Zaregonbadi et al. (2025). These models have been tested in various experimental setups, e.g., Burrage and Sakstein (2016); Pokotilovski (2013a, b); Jenke et al. (2014); Baum et al. (2014); Burrage et al. (2015); Hamilton et al. (2015); Lemmel et al. (2015); Burrage and Copeland (2016); Elder et al. (2016); Ivanov et al. (2016); Burrage et al. (2016); Jaffe et al. (2017); Brax and Pitschmann (2018); Sabulsky et al. (2019); Brax et al. (2018); Cronenberg et al. (2018); Zhang et al. (2019); Arguedas Cuendis et al. (2019); Jenke et al. (2021); Pitschmann (2021); Briddon et al. (2021); Brax et al. (2021); Qvarfort et al. (2022); Yin et al. (2022); Betz et al. (2022); Brax et al. (2023); Zhao et al. (2022); Briddon et al. (2024); Briddon (2023); Fischer et al. (2024a, b); Fischer and Sedmik (2024); Klimchitskaya and Mostepanenko (2024); Haghmoradi et al. (2024); O’Shea et al. (2024); Panda et al. (2024); Pizzuti et al. (2024); Li and Zhu (2025); Sevillano Muñoz (2024); Xiong and Huang (2025); Bachs-Esteban et al. (2025a, b); Altarawneh and Höllwieser (2025); Morales-Navarrete and Cervantes-Cota (2025), while there are also proposals for novel ways of constraining them, see, e.g., Hartley et al. (2019); Burrage et al. (2019b); Käding et al. (2023); Hartley et al. (2024); Tamosiunas et al. (2022a, b); Báez-Camargo et al. (2024); Fischer et al. (2025); Englert et al. (2024); Käding (2025); Lévy and Uzan (2025); Lévy (2024); Garcion et al. (2025); Butt et al. (2025); Nezhadsafavi and Pogosian (2025); Feleppa et al. (2025); Turyshev (2025); Hu et al. (2025); Yuan et al. (2025). The most up-to-date constraints on all three models are provided by Refs. Fischer et al. (2024c); Burrage and Sakstein (2018). However, note that recently it has been suggested that taking into account quantum corrections to symmetrons might significantly weaken their induced fifth forces Udemba and Millington (2025).
@@
Another modified gravity approach is scale-dependent (SD) gravity Reuter and Weyer (2004a); Reuter and Saueressig (2019); Reuter and Weyer (2004b); Percacci (2017); Reuter and Weyer (2004c); Koch and Ramirez (2011); Domazet and Stefancic (2012); Contreras et al. (2016); Koch et al. (2015); Bertini (2025), which is inspired by asymptotic safety Weinberg (1980); Reuter and Saueressig (2019). In this approach, one works with a low-energy effective action that is obtained by integrating out all quantum fluctuations above a certain scale $k(x)$ in the full (unknown) action. The ignorance of the precise UV-structure of the theory then manifests as a running of the gravitational couplings. While the effective action generally consists of infinitely many terms, one usually works with the so-called Einstein-Hilbert truncation, which amounts to standard GR with the gravitational couplings being replaced by functions of $k$. A major advantage of this approach is that it allows to assess quantum properties of gravity without having a concrete UV-completion at hand.
@@
Finally, there is also the theory of $f(R)$-gravity Sotiriou and Faraoni (2010), in which the Ricci scalar in the Einstein-Hilbert action is replaced by an arbitrary function $f(R)$ thereof. For $f(R)$-gravity, it is well-known that it can be embedded into both SD gravity Calzada et al. (2023) and STTs Nojiri et al. (2017). Note that the latter equivalence also holds at the quantum level on-shell, but is broken by off-shell quantum corrections Ruf and Steinwachs (2018); Ohta (2018). However, only specific SD theories and STTs can be expressed as $f(R)$-gravity theories. These relations suggest that there exists also a third equivalence, one between SD gravity and STTs themselves. While special cases of such an equivalence have been encountered in the literature, e.g. in Ref. Reuter and Weyer (2004a), the general case has not been investigated yet. In this work, we will bridge this gap by providing a detailed account of the full equivalence b-etween these two theories. This is done both on the level of the actions and on the level of the field equations. The resulting equivalence relations are simple and can readily be applied to any model satisfying some well-motivated consistency conditions.
@@ II.1 Scalar-tensor theories
STTs of gravity are theories in which additional scalar degrees of freedom are coupled to the metric tensor. While there can be multiple scalar fields in an STT and, generally, the kinetic structure of such models can be rather intricate (see Horndeski Horndeski (1974) and beyond Horndeski theories Gleyzes et al. (2015)), we restrict our discussion to cases with only a single conformally coupled scalar field $\phi$ that has a canonical kinetic term in the so-called Einstein frame, which is one of the two conformal frames prevalently used to formulate STTs in Fujii and Maeda (2007). In this frame, a STT can be described by the action
@@ II.2 Scale-dependent gravity
Here, we will briefly explain the fundamental ingredients and principles of SD gravity. There are many attempts to incorporate quantum effects in the gravitational sector. More precisely, within quantum-inspired gravitational theories, quantum corrections have been consistently incorporated into classical backgrounds, leading to fruitful developments like SD gravity. In general, there are three main approaches to introducing quantum corrections into classical gravitational solutions Reuter and Weyer (2004a); Rincon et al. (2023): at the level of the solutions, at the level of the equations of motion, or at the level of the actions. One of the known and well-understood quantum effects that arises in quantum field theories is the running behavior of the fundamental couplings of the theory. While deriving SD theories from renormalization–group approaches can be questioned Donoghue (2020, 2024), the appearance of effective scale-dependence is a ubiquitous phenomenon. It is expected to arise in a wide range of underlying theories, including the low–energy limit of string theory Hur et al. (2025).
@@
where the Newton constant $G$ and the cosmological constant $\Lambda$ have been replaced by functions depending on the scale-setting relation $k(x)$. The SD equations of motion are obtained by varying the action with respect to the metric. Since the Ricci scalar is coupled to the scale-dependent Newton constant via a term $\frac{\tilde{R}}{G(k)}$, we interpret Eq. (7) as the Jordan frame action of the theory. Note that in practically all SD gravity models $B(k)=0$ is used Rincón (2019); Alvarez et al. (2024a, 2022, b); Rincon et al. (2023). Furthermore, the reader should note that the absence of the kinetic term in the action plays a fundamental role in the formalism of SD gravity, as it enables new families of solutions depending on the functional form of Newton’s coupling. It is also important to emphasize that this absence is critical because, conventionally, it implies that the auxiliary scalar field $k(x)$ is non-dynamical. In the present study, however, by including the kinetic term, we explore a significantly broader sector of the theory and its corresponding solution space. In particular, in the context of an equivalence to STTs, it makes sense to consider a kinetic term for the scale $k(x)$ with an unspecified $B(k)$. Currently, there is no unique way of determining the scale-setting relation $k(x)$. However, progress has been made in finding physically motivated conditions to fix the scale Koch et al. (2015); Alvarez et al. (2024b, a). We will return to this issue in Sec. III.3.
@@ III.1.1 Mapping SD gravity to STTs
When starting from a SD gravity theory, we assume that a scale-setting relation $k(x)$ has been fixed and that the functions $G(k(x))$, $B(k(x))$ and $\Lambda(k(x))$ are known. These quantities can then be used to define $\phi$, $A^{2}(\phi)$ and $V(\phi)$ using the relations (9) - (11), which then determine the STT. Since $G$, $B$ and $\Lambda$ can only depend on $x$ via the scale-setting relation $k(x)$, all expressions in the STT corresponding to these functions must satisfy this requirement as well. From this, we conclude that $\phi(x)\equiv\phi(k(x))$. This is similar to the equivalence of SD gravity and $f(R)$-theory Calzada et al. (2023), where it is necessary to ensure that the function $f$ defined from the SD quantities is indeed only a function of the Ricci scalar.
@@
which is the same scale factor obtained when translating $f(R)$-gravity into a STT Sotiriou and Faraoni (2010); Käding (2019). This does not come as a surprise since we already know that there is an equivalence between $f(R)$-gravity and a SD gravity with a vanishing kinetic term Calzada et al. (2023). This indicates that the equivalence found here complements the existing equivalence relations. We will elaborate more on this in Sec. III.2. For a general $B(k)$, we rewrite Eq. (15) as
@@
Since any well-behaved SD gravity model discussed in the literature satisfies the requirements stated above, see Ref. Rincón (2019) and references therein, we have shown that this class of theories can be fully embedded into STTs. Next, we will see that the situation is different in the opposite direction.
@@ III.2 Consistency check: mapping $f(R)$-gravity onto itself
Figure 1: Triangle depicting the equivalence relations between SD gravity, $f(R)$-gravity and STTs; a solid line means that every model of the original type can be mapped to a theory of the resulting type, while a dotted line means that only specific models of the original type can consistently be mapped. Every SD theory with smooth couplings $G(k)>0$ and $B(k)$ fulfilling Eq. (19) can be uniquely mapped onto a STT. In the other direction, every STT can be mapped onto an SD theory, however, not in a unique way. STTs can only be mapped to $f(R)$-gravity if their scale factors are of a particular form, see Eq. (34), and if Eq. (38) is invertible. For the direction $f(R)$-gravity $\to$ STT, it is required that $f^{\prime}(R)$ is invertible. Ref. Calzada et al. (2023) states that only some SD theories can be mapped to $f(R)$-gravity, while $f(R)$-gravity is always equivalent to SD gravity. Though, the last is in contradiction with our findings, cp. the mappings $f(R)$-gravity $\to$ SD gravity $\to$ STTs and $f(R)$-gravity $\to$ STTs, and we actually conclude from the equations in Ref. Calzada et al. (2023) that the invertibility of $f^{\prime}(R)$ is also required in the direction $f(R)$-gravity $\to$ SD gravity. Therefore, this figure depicts maps from $f(R)$-gravity to SD gravity with a dotted line.
With the equivalence introduced in Sec. III.1 we have completed the equivalence triangle of SD gravity, $f(R)$-gravity and STTs; see Fig. 1. Consequently, we can now perform an important consistency check by mapping $f(R)$-gravity onto itself along the triangle. It is known that $f(R)$-theory can be embedded into SD gravity Calzada et al. (2023) and into STTs Nojiri et al. (2017); Olmo (2007) if $f^{\prime}(R):=df/dR$ is invertible. However, only specific SD theories and STTs can be mapped to an equivalent $f(R)$-theory.
@@
Following Ref. Calzada et al. (2023), we can go from $f(R)$-gravity to a SD theory by defining a cosmological constant
@@
where $G(k(x))$ can be defined arbitrarily [^1] As we have stated in Sec. III.1.1, a defining feature of SD gravity is that the running couplings depend on $x$ only through the scale-setting relation $k(x)$. However, this aspect of the equivalence was not considered in Ref. Calzada et al. (2023), i.e, it was not explained how $k(x)$ needs to be defined for Eq. (28) to be consistently satisfied. We expect that taking this issue into account will lead to additional constraints on either the resulting SD theories or the $f(R)$-theories, for example, the requirement for $f^{\prime}(R)$ to be invertible. Those could enable us to find an explicit expression for $G(k)$, thus uniquely fixing the SD theory. Though, this is beyond the scope of the present paper, which is why we will assume that we can always find a suitable $k(x)$.. As a result, the action in Eq. (27) can be written as
@@
Including STTs into the discussion, we can obtain a STT theory from $f(R)$-gravity via Käding (2019); Olmo (2007)

### v2

@@ I Introduction
Despite the overwhelming success of general relativity (GR), evidence has accumulated over the past century that motivates searches for alternative or modified theories of gravity. In particular, questions about the nature of dark energy and dark matter remain among the most significant problems in gravitational physics and cosmology. While it is possible to account for dark energy within GR in the form of a cosmological constant, this leads to severe fine-tuning problems [1]. In addition, despite substantial theoretical and experimental efforts, the search for direct evidence of particle dark matter has so far been without success. Consequently, modified theories of gravity are extensively discussed in the literature and many different models have emerged [2]. Studying the similarities, differences and interrelations of the various approaches is essential for our understanding of gravity as a whole.
@@
A simple way of modifying GR is by coupling an additional scalar field to the metric tensor, resulting in a so-called scalar-tensor theory (STT) of gravity [3]. This can be motivated, for example, by higher-dimensional theories whose compactified extra dimensions give rise to additional scalar degrees of freedom [4]. Due to the coupling of a scalar field to the metric tensor, we would expect to see an additional fifth force of Nature. However, fifth forces are already heavily constrained by Solar System-based tests [5, 6, 7]. In order to circumvent such experimental bounds, some STTs are equipped with a screening mechanism that renders the fifth force feeble in environments of sufficiently large mass densities. Amongst the most prominent screened scalar field models are the chameleon [8, 9], the symmetron [10, 11, 12, 13, 14, 15, 16, 17], and the environment-dependent dilaton [12, 18, 19, 20, 15, 21, 22]. Chameleons [23] and symmetrons [24, 25, 26], as well as their generalizations [27], have recently been proposed as alternatives to particle dark matter. Screened scalar field models have been tested in various experimental setups. A good overview of such experiments can be found in Ref. [28, 29]. Many tight constraints on screened scalar fifth forces stem from atom interferometry experiments [30, 31, 32, 33, 34, 35, 36], of which Ref. [37] reports the latest results. There are also suggestions of new atom interferometry experiments [38] as well as ideas to improve existing constraints by using Bose-Einstein condensate (BEC) interferometry [39]. Other parts of the model parameter spaces have been constrained by ultra-cold neutron experiments [40, 41, 42, 43, 44], in particular, neutron interferometry [45, 46]. As an extension, neutron Lloyd interferometry was suggested [47, 48]. Additional constraints have been obtained from Lunar Laser Ranging [49], torsion pendulums [50], levitated force sensors [51], hydrogen-like systems [52], and gravitational waves [53]. In particular, solar chameleons are actively searched for [54, 55, 56]. In this context, first measurements have been made with CAST [57]. A recent reassessment of more general Solar System-based tests can be found in [58]. In addition, there have been multiple suggestions for other experiments that could constrain screened scalar fields. Those include mechanical systems [59], quantum optomechanics [60, 61], atomic clocks [62, 63], the LHC [64], terrestrial dark matter direct detection experiments [65], and Casimir force experiments [66, 67]. For the last (and related experiments), there exist analytical studies with exact solutions for the fields [68, 69, 70]. Extensions of Casimir experiments by taking into account the dynamical Casimir effect [71] or quantum and thermal pressures [72, 73] have also been discussed. Accounting for open quantum dynamical effects might, in the future, lead to even tighter bounds on screened scalar fields [74, 75, 76]. In order to optimize experiments searching for screened scalar fields, there are numerical methods [77] and machine learning approaches [78]. In particular, there is SELCIE [79, 80], a tool that allows for studying field profiles around arbitrary sources. In astrophysics and cosmology, there have been suggestions to obtain constraints on screened scalar fields from galaxy clusters [81, 82], gravitational redshift measurements [83] or from non-local quantum correlations between entangled spin pairs [84]. There have a also been studies of screened scalar fields around white dwarfs [85, 86], in NFW halos [87] and in cosmic voids [88]. For symmetrons, in particular, the fullshape galaxy power spectrum [89] and a connection to cosmic strings [90] have been discussed. It has also been suggested to study spacetimes with conformally coupled screened scalar fields in BEC analogue gravity simulations [91], and Ref. [92] provides a particle physics perspective on screening. The most up-to-date constraints on all three models are provided by Refs. [93, 94, 95]. However, note that recently it has been suggested that taking into account quantum corrections to symmetrons might significantly weaken their induced fifth forces [96].
@@
Another modified gravity approach is scale-dependent (SD) gravity which is inspired by asymptotic safety [97, 98]. The scale-dependent formalism in gravity has its roots in several seminal works. A pioneering contribution was made by Reuter and Weyer, who explored the concept of renormalization group-improved gravitational actions [99]. A detailed exposition of the underlying framework, namely the Functional Renormalization Group, can be found in the textbook Quantum Gravity and the Functional Renormalization Group by Reuter and Percacci [98]. The potential implications of quantum Einstein gravity on astrophysical scales are discussed in [100], while the effects of running couplings on galactic rotation curves are analyzed in [101]. A concise introduction to these ideas is provided by Percacci [102]. These developments established the basis for what is now known as scale-dependent gravity. The first explicit application of this framework appeared in [103], with subsequent progress at the cosmological level. Further exploration of the phenomenological consequences of this approach can be found in [104] and [105]. In this approach, one works with a low-energy effective action that is obtained by integrating out all quantum fluctuations above a certain scale $k(x)$ in the full (unknown) action. The ignorance of the precise UV-structure of the theory then manifests as a running of the gravitational couplings. While the effective action generally consists of infinitely many terms, one usually works with the so-called Einstein-Hilbert truncation, which amounts to standard GR with the gravitational couplings being replaced by functions of $k$. A major advantage of this approach is that it allows to assess quantum properties of gravity without having a concrete UV-completion at hand. Note that there have also been other recent discussions on the physical running of the Newton coupling and the cosmological constant [106] with applications to black hole thermodynamics [107] and cosmological models [108, 109].
@@
Finally, there is also the theory of $f(R)$-gravity [110], in which the Ricci scalar in the Einstein-Hilbert action is replaced by an arbitrary function $f(R)$ thereof. For $f(R)$-gravity, it is well-known that it can be embedded into both SD gravity [111] and STTs [112]. Note that the latter equivalence also holds at the quantum level on-shell, but is broken by off-shell quantum corrections [113, 114]. However, only specific SD theories and STTs can be expressed as $f(R)$-gravity theories. These relations suggest that there exists also a third equivalence, one between SD gravity and STTs themselves. While special cases of such an equivalence have been encountered in the literature, e.g. in Ref. [99], the general case has not been investigated yet. In this work, we will bridge this gap by providing a detailed account of the full equivalence b-etween these two theories. This is done both on the level of the actions and on the level of the field equations. The resulting equivalence relations are simple and can readily be applied to any model satisfying some well-motivated consistency conditions.
@@ II.1 Scalar-tensor theories
STTs of gravity are theories in which additional scalar degrees of freedom are coupled to the metric tensor. While there can be multiple scalar fields in an STT and, generally, the kinetic structure of such models can be rather intricate (see Horndeski [115] and beyond Horndeski theories [116]), we restrict our discussion to cases with only a single conformally coupled scalar field $\phi$ that has a canonical kinetic term in the so-called Einstein frame, which is one of the two conformal frames prevalently used to formulate STTs in [3]. Note that the equivalence of conformal frames is still an open question; see, for example, Ref. [117]. In the Einstein frame, a STT can be described by the action
@@ II.2 Scale-dependent gravity
Here, we will briefly explain the fundamental ingredients and principles of SD gravity. There are many attempts to incorporate quantum effects in the gravitational sector. More precisely, within quantum-inspired gravitational theories, quantum corrections have been consistently incorporated into classical backgrounds, leading to fruitful developments like SD gravity. In general, there are three main approaches to introducing quantum corrections into classical gravitational solutions [99, 119]: at the level of the solutions, at the level of the equations of motion, or at the level of the actions. One of the known and well-understood quantum effects that arises in quantum field theories is the running behavior of the fundamental couplings of the theory. While deriving SD theories from renormalization–group approaches can be questioned [120, 121], the appearance of effective scale-dependence is a ubiquitous phenomenon. It is expected to arise in a wide range of underlying theories, including the low–energy limit of string theory [122].
@@
where the Newton constant $G$ and the cosmological constant $\Lambda$ have been replaced by functions depending on the scale-setting relation $k(x)$. The SD equations of motion are obtained by varying the action with respect to the metric. Since the Ricci scalar is coupled to the scale-dependent Newton constant via a term $\frac{\tilde{R}}{G(k)}$, we interpret Eq. (7) as the Jordan frame action of the theory. Note that in practically all SD gravity models $B(k)=0$ is used [123, 124, 125, 126, 119]. Furthermore, the reader should note that the absence of the kinetic term in the action plays a fundamental role in the formalism of SD gravity, as it enables new families of solutions depending on the functional form of Newton’s coupling. It is also important to emphasize that this absence is critical because, conventionally, it implies that the auxiliary scalar field $k(x)$ is non-dynamical. In the present study, however, by including the kinetic term, we explore a significantly broader sector of the theory and its corresponding solution space. In particular, in the context of an equivalence to STTs, it makes sense to consider a kinetic term for the scale $k(x)$ with an unspecified $B(k)$. Currently, there is no unique way of determining the scale-setting relation $k(x)$. However, progress has been made in finding physically motivated conditions to fix the scale [104, 126, 124]. We will return to this issue in Sec. III.3.
@@ III.1.1 Mapping SD gravity to STTs
When starting from a SD gravity theory, we assume that a scale-setting relation $k(x)$ has been fixed and that the functions $G(k(x))$, $B(k(x))$ and $\Lambda(k(x))$ are known. These quantities can then be used to define $\phi$, $A^{2}(\phi)$ and $V(\phi)$ using the relations (9) - (11), which then determine the STT. Since $G$, $B$ and $\Lambda$ can only depend on $x$ via the scale-setting relation $k(x)$, all expressions in the STT corresponding to these functions must satisfy this requirement as well. From this, we conclude that $\phi(x)\equiv\phi(k(x))$. This is similar to the equivalence of SD gravity and $f(R)$-theory [111], where it is necessary to ensure that the function $f$ defined from the SD quantities is indeed only a function of the Ricci scalar.
@@
which is the same scale factor obtained when translating $f(R)$-gravity into a STT [110, 127]. This does not come as a surprise since we already know that there is an equivalence between $f(R)$-gravity and a SD gravity with a vanishing kinetic term [111]. This indicates that the equivalence found here complements the existing equivalence relations. We will elaborate more on this in Sec. III.2. For a general $B(k)$, we rewrite Eq. (15) as
@@
Since any well-behaved SD gravity model discussed in the literature satisfies the requirements stated above, see Ref. [123] and references therein, we have shown that this class of theories can be fully embedded into STTs. Next, we will see that the situation is different in the opposite direction.
@@ III.2 Consistency check: mapping $f(R)$-gravity onto itself
Figure 1: Triangle depicting the equivalence relations between SD gravity, $f(R)$-gravity and STTs; a solid line means that every model of the original type can be mapped to a theory of the resulting type, while a dotted line means that only specific models of the original type can consistently be mapped. Every SD theory with smooth couplings $G(k)>0$ and $B(k)$ fulfilling Eq. (19) can be uniquely mapped onto a STT. In the other direction, every STT can be mapped onto an SD theory, however, not in a unique way. STTs can only be mapped to $f(R)$-gravity if their scale factors are of a particular form, see Eq. (34), and if Eq. (38) is invertible. For the direction $f(R)$-gravity $\to$ STT, it is required that $f^{\prime}(R)$ is invertible. Ref. [111] states that only some SD theories can be mapped to $f(R)$-gravity, while $f(R)$-gravity is always equivalent to SD gravity. Though, the last is in contradiction with our findings, cp. the mappings $f(R)$-gravity $\to$ SD gravity $\to$ STTs and $f(R)$-gravity $\to$ STTs, and we actually conclude from the equations in Ref. [111] that the invertibility of $f^{\prime}(R)$ is also required in the direction $f(R)$-gravity $\to$ SD gravity. Therefore, this figure depicts maps from $f(R)$-gravity to SD gravity with a dotted line.
With the equivalence introduced in Sec. III.1 we have completed the equivalence triangle of SD gravity, $f(R)$-gravity and STTs; see Fig. 1. Consequently, we can now perform an important consistency check by mapping $f(R)$-gravity onto itself along the triangle. It is known that $f(R)$-theory can be embedded into SD gravity [111] and into STTs [112, 128] if $f^{\prime}(R):=df/dR$ is invertible. However, only specific SD theories and STTs can be mapped to an equivalent $f(R)$-theory.
@@
Following Ref. [111], we can go from $f(R)$-gravity to a SD theory by defining a cosmological constant
@@
where $G(k(x))$ can be defined arbitrarily [^1] As we have stated in Sec. III.1.1, a defining feature of SD gravity is that the running couplings depend on $x$ only through the scale-setting relation $k(x)$. However, this aspect of the equivalence was not considered in Ref. [111], i.e, it was not explained how $k(x)$ needs to be defined for Eq. (28) to be consistently satisfied. We expect that taking this issue into account will lead to additional constraints on either the resulting SD theories or the $f(R)$-theories, for example, the requirement for $f^{\prime}(R)$ to be invertible. Those could enable us to find an explicit expression for $G(k)$, thus uniquely fixing the SD theory. Though, this is beyond the scope of the present paper, which is why we will assume that we can always find a suitable $k(x)$.. As a result, the action in Eq. (27) can be written as
@@
Including STTs into the discussion, we can obtain a STT theory from $f(R)$-gravity via [127, 128]